Table 4.2: Enthalpy of Neutralization for an Acid-Base Reaction Trial 1 Concentration of NaOH 1. i mol/L Initial temperatureCAIUTTeter Calorimeter constant or heat capacity of the calorimeter (NOT the specific heat) Average Ccal 31.451°02) Show the required calculations to determine the enthalpy of neutralization (AH) including the Quater and calorimeter
Answer #1

The heat of the calorimeter is calculated:

q cal = Cal * ΔT = 31.4 J / ° C * 6.4 ° C = 200.96 J

Trial 2: 166.42 J

Reaction heat is calculated:

q rx = q water + q cal = 2418 + 200.96 = 2618.96 J

Trial 2: 2000 + 166.42 = 2166.42 J

The moles of reagent limit are 0.05 mol

The reaction enthalpy change is calculated:

ΔHrx = - qrx / n = - 2.62 kJ / 0.05 = - 52.4 kJ / mol

Trial 2: - 43.4 kJ / mol

Average = - 47.9 kJ / mol

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.